A. Find the pressure difference on an airplane wing if air flows over the upper surface with a speed of 113m/s, and along the bottom surface with a speed of 104m/s. Delta P= ? kPa

B. If the area of the wing is 33m^2, what is the net upward force exerted on the wing?

-Express your answer using two sig figs. F= ? kN
Answer #1

From bernoulli:
A =33 m^2
p = 1.29 kg/m^3
Vt = 113 m/s
Vb = 104 m/s

?P = (1/2)*?*(Vt^2-Vb^2)
=1259.685Pa
Lift = ?P * A
= 41569.605 N
Answer #2

From bernoulli's P +1/2 rho v^2 = P + 1/2 rho v^2
so dP = 1/2 * 1.225 ( 113^2 - 104^2) = 1196.2 Pa


b) F=PA = 1196.2*33= 39475 N = 39.475 kN